OpenStudy (anonymous):

5184000 seconds = how many days. Give the exact answer...

OpenStudy (hba):

86,400 minutes

OpenStudy (hba):

1440 hours

OpenStudy (hba):

60 days

OpenStudy (anonymous):

@Devanshi

OpenStudy (anonymous):

this is easy man! Try counting it yourself! A day has 24 into do into 60 seconds! nd divide this from the number you have given!

OpenStudy (hba):

24 hours = 24 x 3,600 seconds = 86,400 seconds = 1 day 1day=86,400 seconds 1*60 days=86,400*60 seconds 60 days=5184000 seconds

OpenStudy (hba):

I cannot make it easier than this :/

OpenStudy (anonymous):

(: sʞuɐɥʇ ʞo